OpenStudy (anonymous):

Find two values of n that satisfy the equation. n(n – 1) = 90 a. 9 and –10 b. 9 and –9 c. 10 and –10 d. 10 and –9 i just need the answer

OpenStudy (anonymous):

n^2 -n-90=0

OpenStudy (anonymous):

to simplify that we need a and b so that ab=90 and a+b=-1

jhonyy9 (jhonyy9):

so a+b=-1 is right but a*b=90 not is because a*b = - 90 hope will help you